OpenStudy (anonymous):

If f(x) = 5x + 3 and g(x) = 2x + 3, what is f(g(x)) ?

OpenStudy (anonymous):

someone help?

OpenStudy (anonymous):

f(x) = 5x + 3 g(x) = 2x + 3 Replacing g(x) in f(x) : f(g(x)) = 5(2x + 3) + 3 =10x + 15 +3 = 10x + 18